OpenStudy (anonymous):

cos(π/∞)=0 ....Can anyone explain by solution.. why?

OpenStudy (anonymous):

there is no such number as \(\frac{\pi}{\infty}\)

OpenStudy (anonymous):

cos is 0 in two places, \[\frac{ \pi }{ 2 }\] and \[\frac{ 3\pi }{ 2 }\]. In the first instance, pi/2 is at 90 degrees reg. We can look at the cos value using a triangle at that angle.. Since cos is defined as x/r, x being 0, and r being 1, you result in your cos value of 0/1, or just 0. The same applies for 3pi/2, only r is -1.